# Scalable UAV Multi-Hop Networking via Multi-Agent Reinforcement Learning with Large Language Models

### v1

@@ Abstract
In disaster scenarios, establishing robust emergency communication networks is critical, and unmanned aerial vehicles (UAVs) offer a promising solution to rapidly restore connectivity. However, organizing UAVs to form multi-hop networks in large-scale dynamic environments presents significant challenges, including limitations in algorithmic scalability and the vast exploration space required for coordinated decision-making. To address these issues, we propose MRLMN, a novel framework that integrates multi-agent reinforcement learning (MARL) and large language models (LLMs) to jointly optimize UAV agents toward achieving optimal networking performance. The framework incorporates a grouping strategy with reward decomposition to enhance algorithmic scalability and balance decision-making across UAVs. In addition, behavioral constraints are applied to selected key UAVs to improve the robustness of the network. Furthermore, the framework integrates LLM agents, leveraging knowledge distillation to transfer their high-level decision-making capabilities to MARL agents. This enhances both the efficiency of exploration and the overall training process. In the distillation module, a Hungarian algorithm-based matching scheme is applied to align the decision outputs of the LLM and MARL agents and define the distillation loss. Extensive simulation results validate the effectiveness of our approach, demonstrating significant improvements in network performance, including enhanced coverage and communication quality.
Index Terms:
Multi-agent reinforcement learning, Large language model, Knowledge distillation, Multi-hop UAV network.
@@ I Introduction
With the rapid advancement of UAV and Internet of Things technologies, extensive research has been devoted to UAV-based communication and networking tasks. Numerous studies have sought to optimize the association and coordination between UAVs and UEs through refined transmission strategies [4], efficient resource management [5], and strategic UAV deployment [6]. Several studies have further explored the applications of UAVs in wireless sensor networks [7], data offloading, and edge computing [8]. Reinforcement learning (RL) techniques, owing to their exceptional decision-making and planning capabilities, have been applied to optimize UAV trajectory planning in dynamic and uncertain environments [9, 8]. These methods aim to improve overall communication performance while addressing the challenges imposed by various operational constraints and environmental uncertainties. However, existing studies primarily focus on the planning of a limited number of UAVs, without addressing the complex coordination required for large-scale multi-UAV systems. In large-scale environments, communication constraints often require the use of multi-hop networks, where connectivity is highly dynamic and unstable. Coordinating UAV swarm within such networks remains a significant challenge.
@@
To address the aforementioned challenges, this paper investigates the application of UAV swarms in multi-hop networking for disaster emergency scenarios, with a focus on optimizing UAV trajectories to maximize both communication coverage and quality under connectivity constraints. To tackle this problem, we propose a novel multi-agent reinforcement learning (MARL) algorithm, referred to as MRLMN (Multi-agent Reinforcement learning with Large language model in Multi-hop Networking). The key contributions of this paper are as follows:
  • This paper formulates the UAV-enabled multi-hop networking task as an optimization problem aimed at maximizing network coverage and communication quality. To address this, a stochastic game framework is adopted to model the dynamic interactions between UAVs and the environment, while capturing the collaboration among UAVs within the network.
  • This paper proposes a task-oriented agent grouping strategy and an information aggregation mechanism within a multi-agent reinforcement learning (MARL) framework. A reward decomposition model is designed based on the grouping strategy to facilitate coordinated decision-making and enhance scalability. Additionally, behavioral constraints are applied to selected UAV groups to improve the robustness and stability of the network.
  • This paper designs a knowledge distillation framework that combines large language models (LLMs) with MARL for UAV networking. A decision matching scheme is proposed to align the outputs of the LLM and MARL agents. Based on this alignment, a tailored distillation loss is designed to transfer high-level decision knowledge, facilitating more effective policy exploration during training.
  • Extensive simulations are conducted to evaluate the proposed approach under various parameters. The results demonstrate its effectiveness and robustness in UAV networking scenarios, showing that the proposed algorithm outperforms existing methods in terms of network coverage and communication quality.
@@ II-A Optimization-Based Methods
In the problem of employing UAVs to provide communication networks for UEs, designing trajectory and resource planning algorithms is complex due to the coupling of various factors. Several studies propose algorithms based on optimization theory to tackle this intricate problem [10, 11]. The majority of these works model the trajectory optimization problem for UAVs within the context of combinatorial optimization problems such as the Traveling Salesman Problem (TSP) and Vehicle Routing Problem (VRP). Building on these well-established frameworks, researchers have developed specialized algorithms to effectively address the optimization challenges present in different UAV-assisted networking scenarios. In [12], block coordinate descent and successive convex optimization techniques are applied to optimize UE communication scheduling and UAV trajectory planning. The work in [13] decomposes the optimization problem into different sub-problems and employs fast global K-means along with an interior-point method to optimize the locations of UAVs. In the context of UAV networking problems, some studies focus on optimizing the uplink and downlink communication rates [12], UAV energy consumption [14, 15], quality of service [16], transmission reliability [17], secrecy communication performance [18] and the number of UEs covered [19]. In [20], a meta-heuristic framework is introduced that employs an artificial bee colony algorithm to solve a mixed-integer programming problem for UAV-assisted communication networks, optimizing UAV 3D positioning under quality of service constraints. Additionally, several studies consider optimizing UAV trajectories in single-cell wireless networks [10], while others take into account environments involving multi-cell scenarios [21]. Under controlled conditions, optimization-based approaches for UAV deployment and planning have demonstrated strong performance and robustness.
@@ II-B RL-Based Methods
In recent years, reinforcement learning (RL) has emerged as a prominent approach for addressing complex optimization challenges in UAV-assisted networks [22, 23], demonstrating remarkable capabilities in sequential decision-making and scheduling tasks. In [24], the authors introduce a constrained deep Q-network to maximize downlink capacity while ensuring comprehensive coverage for all UEs. [25] proposes a dual-attention RL technique to address the time-varying UE traffic and mobility challenges in the environment. Additionally, several studies integrate UAV-assisted networking with data offloading and edge computing. These works employ RL algorithms to achieve fair throughput among UAVs [22, 26] or to minimize the content acquisition delay for UEs [8]. Song et al. [27] propose an improved evolutionary multi-objective RL algorithm to address both trajectory control and task offloading problems of the UAVs.
@@
In this context, MARL [28, 29] algorithms enable each UAV to act as an individual decision-maker while leveraging interactions with other UAVs, making them well-suited for multi-hop wireless networks. [30] utilizes a multi-agent deep deterministic policy gradient (MADDPG) approach to optimize task offloading strategies between UEs and UAVs. Similarly, [31] proposes a decentralized multi-agent soft actor-critic algorithm to optimize spectral efficiency among UAVs. To capture the complex relationships between UAVs and UEs, [9] introduces a heterogeneous graph-based formulation that is integrated into MARL frameworks to facilitate the learning of distributed policies for UAVs. And [32] develops an attention-based heterogeneous graph neural network combined with model-based RL to optimize the UAVs’ resource allocation.
@@
Despite significant advancements in MARL, some issues still persist. Many studies assume that the backhaul network interfacing with the core network is fully configured, thereby neglecting the optimization of relay nodes. In large-scale dynamic environments, such assumptions overlook the scalability challenges of coordinating dozens of UAVs for multi-hop relaying. As the number of UAVs increases, RL techniques encounter convergence issues due to the exponential growth of state and action spaces. The intrinsic randomness of RL further complicates long-term link maintenance in UAV networks. Another critical challenge in MARL is the credit assignment, which complicates the evaluation of each UAV’s contribution when only a global reward signal is available. For instance, the suboptimal performance of a single relay UAV can trigger network disconnections and a rapid decline in the global reward, potentially misleading other agents regarding the efficacy of their policies. Similarly, when overall performance gains are driven by a few critical UAVs, the remaining agents may overestimate their contributions. Moreover, the wide range of environmental states in large-scale scenarios results in a vast exploration space, thereby reducing algorithmic robustness. In the early stages of RL training, random initialization of model parameters leads to counterintuitive decisions that prompt unproductive exploration, ultimately impeding the training process. In response to these challenges, this paper offers targeted solutions.
@@ II-C LLM-Based Methods
The development of LLMs has facilitated their application in robotic planning, where they exhibit strong reasoning and decision-making capabilities [33, 34]. Building on this progress, recent studies have extended the use of LLM agents to UAV planning [35, 36], enhancing the efficiency of the planning process in complex environments. Furthermore, integrating LLMs with RL has emerged as a promising research direction. LLMs have been utilized to shape reward functions [37], enabling more nuanced and context-aware feedback for RL agents. They have also been applied to improve state representation [38] by capturing intricate relationships within the environment, which enhances the agent’s understanding of the environmental states. Additionally, LLMs have been employed to refine action selection [39], providing guidance that allows RL agents to make more informed and strategically aligned decisions, ultimately improving the performance of the system. Despite their potential, the application of LLMs in robotics and UAV planning faces notable challenges. In particular, LLMs’ sensitivity to input prompts can lead to inconsistent decision-making, especially in complex and dynamic tasks that require precise control. Moreover, the gap between the generalized knowledge of LLMs and specific domain requirements further complicates their practical integration, with the issue being more pronounced in problems with complex constraints. Therefore, this paper designs a knowledge distillation module that leverages the strengths of LLMs and MARL to address the multi-UAV networking task.
@@ III-B Communication Model
  • UAV-UE links: For the UAV-UE links, a probabilistic path loss framework is employed to model the distinct characteristics and occurrence probabilities of Line-of-Sight (LoS) and Non-Line-of-Sight (NLoS) conditions. This approach captures the additional path loss caused by environmental factors such as shadowing and scattering, which significantly affect air-to-ground communication in realistic propagation environments. The LoS and NLoS path loss model for UAV $u\in\boldsymbol{\mathcal{U}}$ and UE $m\in\boldsymbol{\mathcal{M}}$ is defined as
    $PL_{u,m}^{\text{LoS}}(t)=20\log(\frac{4\pi f_{c}}{c})+20\log(d_{u,m}(t))+\eta_{\text{LoS}},$ (1)
    $PL_{u,m}^{\text{NLoS}}(t)=20\log(\frac{4\pi f_{c}}{c})+20\log(d_{u,m}(t))+\eta_{\text{NLoS}},$ (2)
@@ III-C Connectivity Constraint
$c_{u}^{\text{UAV}}(t)=\left\{\begin{aligned} 1,&\ \text{if }\exists g\in\boldsymbol{\mathcal{G}},\rho_{g,u}(t)\geq\rho_{\text{th}},\\ &\ \text{or }\exists v\in\boldsymbol{\mathcal{U}}\backslash\{u\},c_{v}^{\text{UAV}}(t)=1,\rho_{u,v}(t)\geq\rho_{\text{th}}\\ 0,&\ \text{otherwise}\end{aligned}\right.$ (11)
@@ III-D Objective Model
$s.t.$ $||\boldsymbol{l}_{u}(t+1)-\boldsymbol{l}_{u}(t)||_{2}\leq\Delta\boldsymbol{l}_{u},\forall u\in\boldsymbol{\mathcal{U}}$ (C1)
@@
where $\kappa$ serves as a weighting factor that balances the objective function. The objective function, as shown in equation (P1), aims to maximize the coverage and data rate of all UEs over the optimization period. The first term represents the number of connected UEs, while the second term captures the data rate achieved by UEs. Constraint (C1) limits the movement of each UAV such that the maximum distance it can travel in a single time slot is $\Delta\boldsymbol{l}_{u}=\omega\Delta t$, where $\omega$ denotes the maximum UAV speed and $\Delta t$ represents the duration of one time slot. Constraint (C2) restricts UAV trajectories to the spatial confines of the environment, denoted by $\boldsymbol{L}^{E}$. Moreover, the optimization problem is further bounded by communication and connectivity constraints, as delineated in equations (11) and (12).
@@ IV-A1 State space $\boldsymbol{\mathcal{S}}$
It defines all possible UAV actions, encompassing both movement direction and speed. This paper assumes that UAVs operate at a fixed speed $\omega$ and their movement is confined to the planar domain, including eight discrete directional movements along with a hovering option.
@@ IV-A3 State transition probability $P$
The reward model $\boldsymbol{\mathcal{R}}$ quantifies UAV actions by evaluating network connectivity and data rate. Accordingly, the team reward, representing the environmental feedback on the effectiveness of UAV networking, is defined as

### v2

@@ Abstract
In disaster scenarios, establishing robust emergency communication networks is critical, and unmanned aerial vehicles (UAVs) offer a promising solution to rapidly restore connectivity. However, organizing UAVs to form multi-hop networks in large-scale dynamic environments presents significant challenges, including limitations in algorithmic scalability and the vast exploration space required for coordinated decision-making. To address these issues, we propose MRLMN, a novel framework that integrates multi-agent reinforcement learning (MARL) and large language models (LLMs) to jointly optimize UAV agents toward achieving optimal networking performance. The framework incorporates a grouping strategy with reward decomposition to enhance algorithmic scalability and balance decision-making across UAVs. In addition, behavioral constraints are applied to selected key UAVs to improve the robustness of the network. Furthermore, the framework integrates LLM agents, leveraging knowledge distillation to transfer their high-level decision-making capabilities to MARL agents. This enhances both the efficiency of exploration and the overall training process. In the distillation module, a Hungarian algorithm-based matching scheme is applied to align the decision outputs of the LLM and MARL agents and define the distillation loss. Extensive simulation results validate the effectiveness of our approach, demonstrating significant improvements in network performance over the MAPPO baseline and other comparison methods, including enhanced coverage and communication quality.
@@ I Introduction
With the rapid advancement of UAV and Internet of Things technologies, extensive research has been devoted to UAV-based communication and networking tasks [5, 49, 7]. Numerous studies have sought to optimize the association and coordination between UAVs and UEs through refined transmission strategies [9], efficient resource management [19], and strategic UAV deployment [6]. Several studies have further explored the applications of UAVs in wireless sensor networks [22, 37], data offloading, and edge computing [18]. Reinforcement learning (RL) techniques, owing to their exceptional decision-making and planning capabilities, have been applied to optimize UAV trajectory planning in dynamic and uncertain environments [52]. These methods aim to improve overall communication performance while addressing the challenges imposed by various operational constraints and environmental uncertainties. However, existing studies primarily focus on the planning of a limited number of UAVs, without addressing the complex coordination required for large-scale multi-UAV systems. In large-scale environments, communication constraints often require the use of multi-hop networks, where connectivity is highly dynamic and unstable. Coordinating UAV swarm within such networks remains a significant challenge.
@@
To address the aforementioned challenges, this paper investigates the application of UAV swarms in multi-hop networking for disaster emergency scenarios, focusing on optimizing UAV trajectories to maximize both communication coverage and quality under connectivity constraints. To tackle this problem, we propose a novel multi-agent reinforcement learning (MARL) framework, referred to as MRLMN (Multi-agent Reinforcement learning with Large language model in Multi-hop Networking). The key contributions of this paper are summarized as follows:
  • This paper formulates the UAV-enabled multi-hop networking task as a multi-objective optimization problem aimed at maximizing network coverage and communication quality while satisfying connectivity constraints. This problem is modeled as a stochastic game to capture the dynamic interactions among UAVs and between UAVs and the environment, explicitly considering collaboration and coordination challenges in large-scale, multi-UAV, and multi-hop network scenarios.
  • This paper proposes a task-oriented agent grouping strategy and an information aggregation mechanism within a MARL framework. A reward decomposition model is designed based on the grouping strategy to facilitate coordinated decision-making, mitigate non-stationarity, and enhance scalability for large UAV swarms. Behavioral constraints are further applied to critical UAV groups to improve network robustness and prevent detrimental topology disruptions.
  • This paper designs a knowledge distillation framework that combines Large Language Models (LLMs) with MARL for UAV networking. In our framework, the LLM serves as an offline advisor that provides high-level strategic insights, which are distilled into the MARL agents to guide policy learning, without being deployed for real-time UAV control. LLM outputs are aligned with MARL actions through a per-agent decision matching scheme. The knowledge is then transferred via a tailored distillation loss, facilitating more efficient policy exploration, addressing the cold-start problem, and enhancing the convergence of MARL training.
  • Extensive simulations are conducted to evaluate the proposed approach across different environment scales, UAV swarm sizes, and baseline comparisons. The results demonstrate that MRLMN consistently outperforms existing methods in terms of network coverage, communication quality, and robustness, validating the effectiveness of the proposed mechanisms.
@@ II Related Work
This section reviews existing research relevant to UAV-assisted networking, covering optimization-based, RL-based, and LLM-assisted methods. The discussion summarizes key developments in each category, identifies their limitations, and emphasizes the advantages of the proposed approach relative to prior work.
@@ II-A Optimization-Based Methods
In the problem of employing UAVs to provide communication networks for UEs, designing trajectory and resource planning algorithms is complex due to the coupling of various factors. Several studies propose algorithms based on optimization theory to tackle this intricate problem [32]. The majority of these works model the trajectory optimization problem for UAVs within the context of combinatorial optimization problems such as the Traveling Salesman Problem (TSP) and Vehicle Routing Problem (VRP). Building on these well-established frameworks, researchers have developed specialized algorithms to effectively address the optimization challenges present in different UAV-assisted networking scenarios. In [42], block coordinate descent and successive convex optimization techniques are applied to optimize UE communication scheduling and UAV trajectory planning. The work in [12] decomposes the optimization problem into different sub-problems and employs fast global K-means along with an interior-point method to optimize the locations of UAVs. In the context of UAV networking problems, some studies focus on optimizing the uplink and downlink communication rates [42], UAV energy consumption [35], quality of service [46] and the number of UEs covered [27]. Under controlled conditions, optimization-based approaches for UAV deployment and planning have demonstrated strong performance and robustness.
@@
To address these limitations, the proposed approach leverages MARL to efficiently explore high-dimensional state and action spaces, reducing reliance on precise modeling of the networking task. Task-oriented grouping and reward decomposition simplify the coordination among multiple UAVs, mitigating the computational complexity associated with large-scale deployments. These combined mechanisms allow the framework to generate effective and robust UAV deployment decisions in dynamic, large-scale multi-hop networking scenarios.
@@ II-B RL-Based Methods
In recent years, reinforcement learning (RL) has emerged as a prominent approach for addressing complex optimization challenges in UAV-assisted networks [33], demonstrating remarkable capabilities in sequential decision-making and scheduling tasks. In [51], the authors introduce a constrained deep Q-network to maximize downlink capacity while ensuring comprehensive coverage for all UEs. [3] proposes a dual-attention RL technique to address the time-varying UE traffic and mobility challenges in the environment. Additionally, several studies integrate UAV-assisted networking with data offloading and edge computing. These works employ RL algorithms to achieve fair throughput among UAVs [38] or to minimize the content acquisition delay for UEs [17]. Song et al. [29] propose an improved evolutionary multi-objective RL algorithm to address both trajectory control and task offloading problems of the UAVs.
@@
In this context, MARL [8] algorithms enable each UAV to act as an individual decision-maker while leveraging interactions with other UAVs, making them well-suited for multi-hop wireless networks. [14] utilizes a multi-agent deep deterministic policy gradient (MADDPG) approach to optimize task offloading strategies between UEs and UAVs. Similarly, [25] proposes a decentralized multi-agent soft actor-critic algorithm to optimize spectral efficiency among UAVs. To capture the complex relationships between UAVs and UEs, [52] introduces a heterogeneous graph-based formulation that is integrated into MARL frameworks to facilitate the learning of distributed policies for UAVs. And [26] develops an attention-based heterogeneous graph neural network combined with model-based RL to optimize the UAVs’ resource allocation.
@@
Despite significant advancements in MARL, some issues still persist. Many studies assume that the backhaul network interfacing with the core network is fully configured, thereby neglecting the optimization of relay nodes. In large-scale dynamic environments, such assumptions overlook the scalability challenges of coordinating dozens of UAVs for multi-hop relaying. As the number of UAVs increases, RL techniques encounter convergence issues due to the exponential growth of state and action spaces. The intrinsic randomness of RL further complicates long-term link maintenance in UAV networks. Another critical challenge in MARL is the credit assignment, which complicates the evaluation of each UAV’s contribution when only a global reward signal is available. For instance, the suboptimal performance of a single relay UAV can trigger network disconnections and a rapid decline in the global reward, potentially misleading other agents regarding the efficacy of their policies. Similarly, when overall performance gains are driven by a few critical UAVs, the remaining agents may overestimate their contributions. Moreover, the wide range of environmental states in large-scale scenarios results in a vast exploration space, thereby reducing algorithmic robustness. In the early stages of RL training, random initialization of model parameters leads to counterintuitive decisions that prompt unproductive exploration, ultimately impeding the training process.
In response to these challenges, in the proposed framework, UAVs are coordinated through a task-oriented grouping strategy that assigns specific roles, ensuring that relay responsibilities are explicitly considered. To mitigate the scalability and convergence issues arising from high-dimensional state and action spaces, a reward decomposition scheme is applied, which distributes global feedback into more localized signals, thereby improving credit assignment for individual UAVs and stabilizing learning. Behavioral constraints are enforced on critical UAVs to prevent disruptive topology changes, ensuring network robustness despite the intrinsic randomness of RL. Finally, to reduce the burden of exploration in vast and complex environments, large language models are employed during offline training to provide high-level strategic guidance, which is distilled into MARL policies to accelerate learning, guide effective UAV deployment, and maintain decentralized decision-making suitable for real-time operation.
@@ II-C LLM-Based Methods
The development of LLMs has facilitated their application in robotic planning, where they exhibit strong reasoning and decision-making capabilities [36, 54]. Building on this progress, recent studies have extended the use of LLM agents to UAV planning [43, 44], enhancing the efficiency of the planning process in complex environments. Furthermore, integrating LLMs with RL has emerged as a promising research direction. LLMs have been utilized to shape reward functions [11], enabling more nuanced and context-aware feedback for RL agents. They have also been applied to improve state representation [34] by capturing intricate relationships within the environment, which enhances the agent’s understanding of the environmental states. Additionally, LLMs have been employed to refine action selection [39], providing guidance that allows RL agents to make more informed and strategically aligned decisions, ultimately improving the performance of the system. Despite their potential, the application of LLMs in robotics and UAV planning faces notable challenges. In particular, LLMs’ sensitivity to input prompts can lead to inconsistent decision-making, especially in complex and dynamic tasks that require precise control. Moreover, the gap between the generalized knowledge of LLMs and specific domain requirements further complicates their practical integration, with the issue being more pronounced in problems with complex constraints. Existing methods that integrate LLMs with RL often assume low-dimensional or structured action and state spaces and rely on direct translation of LLM outputs into agent actions or hierarchical abstractions. These assumptions are unsuitable for multi-UAV multi-hop networking, where action and state spaces are high-dimensional, strongly coupled, and constrained by connectivity and communication requirements.
In the proposed approach, the semantic-level reasoning of the LLM is combined with the robustness and coordination capability of MARL to address these challenges. During MARL training, the LLM provides high-level strategic guidance for UAV deployment and networking. MRLMN then aligns these strategies with decentralized agent behaviors by assigning LLM-suggested roles via the Hungarian algorithm and distilling the resulting priors into the MARL policies. At deployment, MARL agents execute decisions independently, ensuring scalable and distributed control without requiring online interaction with the LLM.
@@ III System Model
In this section, the system components of the UAV-enabled emergency network and their spatial dynamics within the disaster environment are defined. The communication model is then introduced together with the connectivity constraints required to maintain a feasible multi-hop backhaul link to the core network. Building on these elements, the networking task is formulated as a joint trajectory optimization problem that aims to maximize user coverage and communication quality under practical operational constraints.
@@ III-B Communication Model
[…]
@@ III-C Connectivity Constraint
[…]
@@ III-D Objective Model
[…]
@@
where $\kappa$ serves as a weighting factor that balances the objective function. The objective function, as shown in equation (P1), aims to maximize the coverage and data rate of all UEs over the optimization period. The first term represents the number of connected UEs, while the second term captures the data rate achieved by UEs. Constraint (C1) limits the movement of each UAV such that the distance it can travel in a single time slot is $\Delta\boldsymbol{l}_{u}=\omega\Delta t$, where $\omega$ denotes the UAV speed and $\Delta t$ represents the duration of one time slot. Constraint (C2) restricts UAV trajectories to the spatial confines of the environment, denoted by $\boldsymbol{L}^{E}$. Moreover, the optimization problem is further bounded by communication and connectivity constraints, as delineated in equations (11) and (12).
@@ IV Methodology
To address the scalability challenges and extensive exploration space in large-scale UAV networking, this section introduces the MRLMN framework. The networking objective is reformulated as a stochastic game, and a decentralized learning architecture based on the IPPO is adopted. Coordination is facilitated through an information aggregation mechanism and a role-based grouping strategy with decomposed rewards, while behavioral constraints ensure feasible motion and communication. A knowledge distillation module is employed to integrate LLM guidance, significantly improving exploration efficiency and guiding the learning of agent policies in complex decision scenarios.
@@ IV-A1 State space $\boldsymbol{\mathcal{S}}$
The action space specifies the set of movement decisions available to each UAV. In this paper, the UAVs move at a fixed cruise speed $\omega$, which is treated as a constant system parameter. Accordingly, the action space includes only the selection of movement direction within the planar domain, modeled as eight discrete directional options together with a hovering action. This formulation captures the essential trajectory control while simplifying the decision-making complexity for MARL agents.
@@ IV-A3 State transition probability $P$
[…]